OpenStudy (anonymous):

44 out of 80 is what %

OpenStudy (anonymous):

55%

OpenStudy (anonymous):

how did u get 55%

OpenStudy (anonymous):

44/80=11/20. Multiply ANY number by 100 and that will give you the %. So (11/20)* 100=55%

OpenStudy (anonymous):

46 is 92% of what number?

OpenStudy (anonymous):

46= (92/100)*x. x= 4600/92=50